### v1

# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

Abstract
In real-world scenarios, scanned point clouds are often incomplete due to occlusion issues. The task of self-supervised point cloud completion involves reconstructing missing regions of these incomplete objects without the supervision of complete ground truth. Current self-supervised methods either rely on multiple views of partial observations for supervision or overlook the intrinsic geometric similarity that can be identified and utilized from the given partial point clouds. In this paper, we propose MAL-SPC, a framework that effectively leverages both object-level and category-specific geometric similarities to complete missing structures. Our MAL-SPC does not require any 3D complete supervision and only necessitates a single partial point cloud for each object. Specifically, we first introduce a Pattern Retrieval Network to retrieve similar position and curvature patterns between the partial input and the predicted shape, then leverage these similarities to densify and refine the reconstructed results. Additionally, we render the reconstructed complete shape into multi-view depth maps and design an adversarial learning module to learn the geometry of the target shape from category-specific single-view depth images. To achieve anisotropic rendering, we design a density-aware radius estimation algorithm to improve the quality of the rendered images. Our MAL-SPC yields the best results compared to current state-of-the-art methods. We will make the source code publicly available at https://github.com/ltwu6/malspc.
Index Terms:
Point cloud completion, shape completion, self-supervised learning, generative adversarial learning
I Introduction
Point clouds, which represent 3D instances as a collection of points in 3D space, have attracted increasing interest due to the growing popularity of 3D scanning devices such as LiDAR and depth cameras. Point clouds offer a detailed representation of the geometric features and spatial structures of real-world 3D objects and scenes. They are widely applied in various fields, such as autonomous driving[1, 2, 3, 4, 5], motion analysis[6, 7], and augmented reality[8, 9]. Accurate and complete shape representation plays a vital role in understanding 3D point clouds. However, point clouds scanned in practical scenarios are often incomplete due to occlusion. This incompleteness severely impacts the performance of downstream tasks, such as 3D object detection [10, 11, 12] and scene understanding [13, 14].
To solve this problem, point cloud completion methods have emerged. Point cloud completion aims to infer the missing parts of an object based on the incomplete point cloud. Most of the existing point cloud completion methods rely on a large number of complete point clouds as supervision [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, in real-world scenes, clean and complete point clouds are really few since current devices struggle to directly capture such data.
To overcome these limitations, some research has focused on unsupervised learning approaches for point cloud completion. Among these methods, some employ complete shapes unpaired with the partial input as priors to drive the completion process [25, 26, 27, 28, 29, 30, 31], while others utilize 2D signals such as depth maps or silhouette images from multiple viewpoints as supervision [32, 33]. However, the former still requires complete point clouds to acquire shape priors, and the latter relies on 2D images of objects and the corresponding projection parameters. Recently, self-supervised methods, which only require partial inputs for supervision, have been proposed. These methods estimate the missing parts of the point cloud by utilizing the multi-view consistency of partial scans and the robust learning capabilities of the masked models. However, they face some limitations. First, these methods rely on multi-view partial point clouds of each object as shape constraints or as training data to drive the shape learning process. Nevertheless, obtaining point clouds from multiple viewpoints is also impractical in real-world applications. Secondly, they fail to utilize the inherent local similarities and category-specific geometric properties of objects.
From our perspective, it is commonly observed that an object frequently exhibits similar or repetitive local structures. For instance, a table typically features four identical cylindrical legs. This observation implies that the existing and missing sections of a point cloud are likely to share similar local geometric patterns. Moreover, objects within the same category generally display consistent structures. For example, most airplanes feature two flat wings. Consequently, these inherent geometric similarities, specific to each category, can be utilized to guide the completion of missing regions in point clouds.
In this paper, we propose MAL-SPC, a Multi-view Adversarial Learning framework for Self-supervised Point cloud Completion. One of the core components, the Pattern Retrieval Network (PRN), reconstructs complete shapes by initially generating a coarse shape and subsequently encoding local geometric patterns from both the coarse and partial input shapes. These patterns manifest in two forms: position encodings, quantified as the average distance between each point and its nearest neighbors, and curvature encodings, defined by the variations in normals between each point and its closest neighbors. Subsequently, a cross-attention-based module is introduced to retrieve correlated local patterns for the coarse shape from the partial shape. These retrieved patterns are further utilized to guide the densification and refinement of the coarse shape, capitalizing on the object-level geometric similarities and repetitiveness inherent in the data.
In addition, we introduce a Multi-view Adversarial Network (MAN) to learn the shared geometry of each category. Objects within the same category commonly exhibit remarkably similar shape distributions because they share many typical geometric structures. This similarity suggests that missing parts of a point cloud are likely to resemble the geometric structure of some objects in the same category. To exploit this similarity, we employ an adversarial learning strategy to learn the distinctive shape characteristics from category-specific data. Specifically, since depth maps provide a more compact and regular shape representation than partial point clouds and are easier to acquire, we render the reconstructed complete point cloud into depth maps from multiple viewpoints to capture the view-specific structures of the shape. Then, we utilize a 2D CNN-based network to discriminate the rendered depth maps and those from the same category in the dataset, thus refining our predictions based on category-specific geometric knowledge. Additionally, to address the anisotropic rendering issue, we propose a density-aware radius estimation algorithm to adaptively adjust the point radius from different viewpoints, thereby improving the quality of rendered images.
Our MAL-SPC is unique in that it does not require multiple views of a partial point cloud for supervision or as additional data to drive the training process. Instead, it can achieve remarkable performance by employing only a single partial point cloud per object for training and supervision. Additionally, MAL-SPC effectively leverages the intrinsic geometry similarity from partial data to reconstruct missing structures. Moreover, unlike other self-supervised methods, our approach does not require any data preprocessing processes, such as random region removal, which is time-consuming and computationally intensive.
We conduct extensive experiments on both synthetic datasets and real-world scans to demonstrate that our model significantly outperforms the state-of-the-art approaches by a large margin.
In summary, the main contributions of this paper are two-fold:
  • We propose a Pattern Retrieval Network that leverages the intrinsic similarity and repetitiveness of object-level local geometry to reconstruct complete shapes. This network effectively infers the missing structures by learning correlated geometric patterns, including position and curvature, from the partial shape.
  • We propose a Multi-view Adversarial Network that facilitates the completion of missing shapes by leveraging the category-specific geometric similarities. To achieve anisotropic rendering and improve the quality of the rendered depth maps, we design a density-aware radius estimation algorithm.
The remainder of this paper is organized as follows. Section II briefly reviews existing supervised and unsupervised techniques for point cloud completion and reconstruction. Section III presents a detailed description of our proposed methods. Section IV discusses our experimental design, findings, and analysis. Finally, Section V concludes this paper.
Figure 1: Flowchart of the proposed self-supervised point cloud completion method. Given a depth map $\mathbf{D}_{0}$, we can obtain an incomplete point cloud $\mathbf{P}_{in}$ by back-projecting it into 3D space. We first employ an auto-encoder to generate a coarse shape, and then introduce a pattern retrieval module to guide the refinement and densification of the shape based on the position and curvature patterns of $\mathbf{P}_{in}$. Then, we render the generated point clouds into silhouette maps from the input viewpoint and depth maps from various random viewpoints. The silhouette maps are supervised by the mask image binarized from $\mathbf{D}_{0}$, while the depth maps are used for adversarial learning with the counterparts from the same category.
II Related Work
II-A Supervised Point Cloud Completion
The supervised point cloud completion methods utilize complete and clean point clouds paired with partial inputs to guide and drive the completion process. Based on the data representation in the encoding and decoding process, the point cloud completion methods can be classified into voxel-based and point-based approaches.
The voxel-based methods [34, 35, 36, 37, 38, 39] first convert partial point clouds into 3D grids through voxelization and subsequently utilize the 3D Convolutional Neural Networks for learning the complete shapes. Dai et al. [34] employed a 3D-Encoder Predictor Network to generate an initial approximation of a shape, followed by a 3D shape synthesis method that refines the coarse output using a shape prior derived from a shape database. Xie et al. [35] introduced a Gridding module along with a reverse module for seamless conversion between point clouds and voxels, thereby minimizing structural information loss during the voxelization process. Additionally, they devised a Cubic Feature Sampling module to extract contextual features from neighboring points. Sun et al. [36] represented point clouds as occupancy grids and leveraged local patches of partial inputs to infer missing parts by capitalizing on the potential repetition of local patterns within an object. Initially, they employed a network to capture similar local patches of the inputs. Then, these patches undergo deformation and blending processes, ultimately reconstructing into the final outputs. Voxel-based methods can utilize the powerful feature extraction capabilities of CNNs. However, the computation speed of 3D CNN is relatively slow, and the conversion process between point clouds and voxels inevitably introduces accuracy loss.
The point-based methods [15, 40, 41, 42, 43, 44, 45] directly encode and decode features within the 3D point space by using Multi-Layer Perceptron (MLP). To obtain smooth and detailed surfaces, some efforts employed the Folding operation, originally proposed in FoldingNet [40], to generate complete point clouds. Yuan et al. [15] designed an MLP-based encoder to extract the global feature based on the partial input. Then, they leveraged fully connected (FC) layers to reconstruct a coarse point cloud. Subsequently, the Folding operation was applied to each coarse point. This approach effectively capitalizes on the inherent flexibility of FC layers and the desirable surface-smoothing characteristics offered by the Folding operation. Groueix et al. [46] employed multiple MLPs to deform 2D grids into distinct local patches, enabling the generation of more intricate shapes compared to FoldingNet. However, this approach resulted in overlapping regions between patches. To address this issue, Liu et al. [47] proposed an expansion penalty method to minimize the overlap areas. Besides, it is difficult for a fixed 2D grid to capture the intricate topologies of point clouds. To improve the shape representation ability of the seeds, Yang et al. [48] design a model for learning high-dimensional seeds for generating complete shapes. Likewise, Pang et al. [49] utilized a network to learn the point cloud topology and tore the grid into patches to better represent the object’s shape.
To achieve more flexible generation of missing points, some studies directly utilize MLPs to generate the 3D positions of points [41, 18, 50, 16, 51, 52, 43, 17, 42, 53]. Inspired by the process of tree growth, Tchapmi et al. [41] proposed a rooted-tree decoder that splits points layer by layer. Xiang et al. [18] introduced a coarse-to-fine approach where each parent point is divided into multiple child points. To drive the splitting process, they designed a skip-transformer block to capture the relationship between the features of child points and parent points. Drawing upon the remarkable capabilities of the Transformer [50] architecture in the fields of image processing and natural language processing (NLP), several recent advancements have introduced point cloud completion methods with Transformer as the backbone[54, 55]. Yu et al. [16] first considered the process of inferring missing points based on the partial points as a translation task. They proposed a Transformer architecture for encoding the points’ features and generating the missing parts of the inputs. Zhou et al. [52] introduced Patch Seeds, which can preserve global structures and regional details to represent the shape. Based on the extracted seed features, they designed an Upsample Transformer to generate points. Chen et al. [17] proposed to use a set of learned anchors to represent the regional information of objects. Then, they incorporated a modulation scheme and 2D grids to detail the sparse points. Li et al. [42] leveraged proxies to represent the existing and missing parts of the shapes and devised a missing part sensitive transformer to refine the missing proxies and thus generate high-quality missing structures. Zhu et al. [53] first proposed a network that utilizes multi-view depth images to reconstruct a coarse shape. Then, they leveraged shape priors and geometric self-similarities to detail and refine the shapes.
In addition to partial point clouds, some methods incorporate view images, including RGB images and depth maps, to assist in point cloud completion [56, 57, 58, 59, 60]. Zhang et al. [57] leveraged a single-view RGB image to reconstruct a coarse point cloud and then devised an offset predictor model to refine it. Zhu et al. [58] predicted the coarse shape by transferring the image shape feature to the 3D point cloud domain and then proposed a dual-refinement module to refine the coarse shape. Hu et al. [59] rendered depth maps of a partial point cloud from different viewpoints and completed the shape by performing depth image completion for each view. To further constrain the geometric consistency among multi-view depth maps of a shape, Hu et al. [60] proposed a consistency loss to preserve the similarity of the shapes generated by multi-view depth maps.
II-B Unsupervised Point Cloud Completion
Due to the challenges in acquiring complete point clouds in real-world scenarios, some studies have proposed unsupervised methods for point cloud completion.
Unsupervised point cloud methods do not require paired complete 3D ground truth to drive the completion process. These methods can be divided into three categories according to the supervision signal, i.e., unpaired methods, 2D-supervised methods, and self-supervised methods. The unpaired point cloud completion methods generate complete point clouds without paired ground truth as supervision during the training process [25, 26, 27, 28, 29, 30, 31, 29, 61]. In fact, these methods leverage complete and clean point clouds from other domains to train a model to capture 3D shape priors. Then, they used the learned shape priors to guide the completion of the partial point cloud from the source domain. The first unpaired method is PCL2PCL, proposed by Chen et al. [25]. PCL2PCL first used partial point clouds from the source domain to pre-train a partial-shape reconstruction Auto-encoder and then trained a complete-shape reconstruction Auto-encoder by utilizing complete point clouds from another domain. Subsequently, they employed a GAN-based model [62] to transform the partial global features to the same distribution with the complete shapes’ global features. To enhance the learning ability of the generation network, Wen et al. [26] proposed a cycle-based model to learn the transformation between partial and complete point clouds. Cao et al. [29] proposed a knowledge transfer model to learn the shape information of partial inputs. They designed a teacher-assistant-student network to transfer the knowledge learned from complete shapes to partial shapes. In addition to the shape priors provided by complete point clouds, Ma et al. [30] proposed to leverage the symmetry of objects to reconstruct the missing regions.
The 2D-supervised methods employ 2D images along with projection parameters to guide the completion of 3D shapes. Wu et al. [32] projected reconstructed point clouds into the 2D plane and then used 2D points within the silhouette maps to supervise the view-specific shape. They also proposed a view-based calibrator to reduce noises in the predicted complete shapes. Aiello et al. [33] enhanced the 3D point clouds’ features with 2D RGB images by attention modules. They introduced a framework that rendered predicted point clouds from certain view angles and used 2D silhouette maps to supervise the renderer images.
The self-supervised methods only leverage the partial point cloud as a supervision signal to guide the completion process. Mittal et al. [63] proposed the first learning-based self-supervised method, PointPnCNet. PointPnCNet first randomly removed a small region from the input partial point cloud and used a network to predict the missing regions, including the manually removed parts and naturally missing parts. It employed the partial matching loss to preserve the existing parts in the predicted shapes. Inspired by closed-loop systems in control theory, Hong et al. [64] first predicted a coarse point cloud and used it to synthesize partial point clouds to input to the same model as the initial input. By utilizing consistency shape loss among predictions generated by multiple partial inputs, the model could generate complete shapes. Cui et al. [65] divided partial point clouds into three groups and then proposed a masked autoencoder to reconstruct complete shapes. Rather than directly using partial Chamfer Distance loss, they designed a Region-aware Chamfer Distance loss to reduce the noises of the predictions. Kim et al. [66] proposed a pose-aware approach for self-supervised point cloud completion. They projected the shape features onto various viewpoints and reconstructed the corresponding incomplete point clouds. They leveraged the shape consistency among multiple incomplete point clouds of the same object to constrain the reconstruction process.
III Proposed Method
Let $\mathbf{D}_{0}\in\mathbb{R}^{H\times W}$ be the depth map scanned from an object. After back-projection, $\mathbf{D}_{0}$ yields a partial point cloud, denoted as $\mathbf{P}_{in}\in\mathbb{R}^{N\times 3}$. $\mathbf{P}_{com}\in\mathbb{R}^{M\times 3}$ represents the complete shape of $\mathbf{P}_{in}$, where $H$ and $W$ denote the height and width of the depth image, $N$ and $M$ are the numbers of points. Our goal is to find a mapping $f(\mathbf{P}_{in})=\mathbf{P}_{out}$ to make the predicted complete shape $\mathbf{P}_{out}$ as close as possible to $\mathbf{P}_{com}$ without $\mathbf{P}_{com}$ as supervision.
As depicted in Figure 1, our framework, namely MAL-SPC, first leverages a Pattern Retrieval Network to generate complete point clouds. Then, we render the complete predictions into a silhouette map and multi-view depth maps. The rendered depth maps are used to perform adversarial learning with the counterparts from the same categories. In what follows, we will provide a detailed explanation of our approach.
III-A Pattern Retrieval Network
Since local geometric patterns of most objects exhibit similarities and repetitions, the missing regions of an object are likely to share similar or even identical local geometric patterns to the existing parts. To leverage this object-level geometric similarity, we propose a cross-attention-based pattern retrieval network to infer the missing structures.
As shown in Figure 2, consuming as input a partial point cloud $\mathbf{P}_{in}$, we first use a 3D encoder derived from PCN [15] to learn a global feature $\mathbf{o}$ that encapsulates the object’s shape information. Then we employ three layers of fully-connected layers to reconstruct a coarse yet complete point cloud $\mathbf{P}_{c}\in\mathbb{R}^{N\times 3}$. Next, we extract the position encodings and curvature encodings from the partial input and coarse output, respectively. The position encodings represent the positional relationship of local points, which are formulated as the average distance between each point and its nearest neighbors. Specifically, for each point $\mathbf{p}_{i}\in\mathbf{P}_{in}$, we first find out its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})=\{\mathbf{q}_{i}^{j}\}_{j=1}^{K}\subset\mathbf{P}_{in}$ and then compute the average Euclidean distance between $\mathbf{p}_{i}$ and $\mathcal{N}_{K}(\mathbf{p}_{i})$, i.e.,
$\mathbf{f}_{pos}^{in}(i)=\frac{1}{K}\sum_{j=1}^{K}||\mathbf{p}_{i}-\mathbf{q}_{i}^{j}||_{2}.$ (1)
The curvature encodings capture the curvature variations across the surfaces of the point cloud. We employ the normals of individual points to characterize surface curvature. For each point $\mathbf{p}_{i}$, we first compute the covariance matrix $\mathbf{C}_{i}$ based on its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})$ determined previously. The covariance matrix captures the local surface geometry and is defined as follows:
$\mathbf{C}_{i}=\frac{1}{K}\sum_{\mathbf{q}_{i}^{j}\in\mathcal{N}_{K}(\mathbf{p}_{i})}(\mathbf{q}_{i}^{j}-\hat{\mathbf{p}_{i}})(\mathbf{q}_{i}^{j}-\hat{\mathbf{p}}_{i})^{T},\hat{\mathbf{p}}_{i}=\frac{1}{K}\sum_{j=1}^{K}\mathbf{q}_{i}^{j}.$ (2)
Figure 2: The flowchart of our Pattern Retrieval Network. “FC” and “MLP” denote the Fully-Connected layer and the Multi-Layer Perceptron. Given a partial point cloud $\mathbf{P}_{in}$, we first utilize an Auto-encoder to generate a coarse shape $\mathbf{P}_{c}$. Next, we separately extract position encodings and curvature encodings for both $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$, and then leverage a cross-attention-based module to identify and retrieve the most related encodings for each point in $\mathbf{P}_{c}$. Subsequently, $\mathbf{P}_{c}$ and the selected encodings are upsampled and fused to refine the shape.
The estimated normal vector $\mathbf{n}_{i}$ for each point $\mathbf{p}_{i}$ is obtained by computing the eigenvector corresponding to the smallest eigenvalue of the covariance matrix $\mathbf{C}_{i}$. Besides, [67] is employed to perform the sign disambiguation of the normals.
To simulate the curvature variations of the surface, inspired by [65], we calculated the cosine similarity between the normal vectors of each point and its neighboring $K$ points.
$v(i,j)=1-\texttt{cos}(\mathbf{n}_{i},\mathbf{n}_{j})=1-\frac{\mathbf{n}_{i}\cdot\mathbf{n}_{j}}{\texttt{max}(||\mathbf{n}_{i}||\cdot||\mathbf{n}_{j}||,\epsilon)},$ (3)
$\mathbf{f}_{cur}^{in}(i)=\sqrt{\frac{1}{K}\sum_{j=1}^{K}(v(i,j)-\frac{1}{K}\sum_{j=1}^{K}v(i,j))^{2}},$ (4)
where $\epsilon$ is a constant to prevent the denominator from being 0.
In this way, we obtain the position encoding $\mathbf{f}_{pos}^{in}\in\mathbb{R}^{N\times 1}$ and curvature encoding $\mathbf{f}_{cur}^{in}\in\mathbb{R}^{N\times 1}$ from the partial input. Similarly, we compute the position encoding $\mathbf{f}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and curvature encoding $\mathbf{f}_{cur}^{c}\in\mathbb{R}^{N\times 1}$ from the coarse shape using the same method. Following this, we introduce a cross-attention-based pattern retrieval module to capture correlated local patterns for coarse predictions from partial inputs. Specifically, for the position encoding, we utilize MLPs to project $\mathbf{f}_{pos}^{in}$ and $\mathbf{f}_{pos}^{c}$ to get queries $\mathbf{g}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and keys $\mathbf{h}_{pos}^{in}\in\mathbb{R}^{N\times 1}$, respectively. Then we employ the algorithm in cross-attention [50] to calculate the weight maps between $\mathbf{h}_{pos}^{in}$ and $\mathbf{g}_{pos}^{c}$, representing the correlation of the position encoding for each point in $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$:
$\mathbf{W}_{pos}=\texttt{Softmax}(\mathbf{h}_{pos}^{in}\cdot\mathbf{g}_{pos}^{c}).$ (5)
Figure 3: Depth maps rendered from point clouds with an equal point number. The rendering process employs a fixed and identical point radius. It is obvious that, with a small foreground area, the visible surface can be well represented. However, when the foreground area is large, a fixed point radius cannot cover the entire visible surface, resulting in some foreground pixels being covered by invisible points.
Next, based on $\mathbf{W}_{pos}$, we retrieve the $L$ most correlated position encodings from $\mathbf{f}_{pos}^{in}$ for each point in $\mathbf{P}_{c}$, denoted as $\mathbf{e}_{pos}\in\mathbb{R}^{N\times L}$. Similarly,the related curvature encodings $\mathbf{e}_{cur}\in\mathbb{R}^{N\times L}$ for each point in $\mathbf{P}_{c}$ are computed in the same manner. Then we upsample $\mathbf{P}_{c}$ along with $\mathbf{e}_{pos}$, $\mathbf{e}_{cur}$, $\mathbf{f}_{pos}^{c}$, $\mathbf{f}_{cur}^{c}$, $\mathbf{o}$, and fuse them in a concatenation manner. Finally, we employ MLPs to infer the point-wised offsets for the upsampled $\mathbf{P}_{c}$ and add them to obtain the final outputs $\mathbf{P}_{out}$.
III-B Multi-view Adversarial Learning
Objects of the same category share many similar structures, and therefore, their shape distributions are close. To make the generated complete point cloud as realistic and reasonable as possible, we develop an adversarial learning module to learn the category-specific shape knowledge from the real depth maps. Specifically, we first employ a silhouette renderer to render $\mathbf{P}_{out}$ and $\mathbf{P}_{c}$ to two silhouette maps $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ from the same viewpoint as $\mathbf{D}_{0}$. $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ are supervised by $\mathbf{D}_{0}$. Then we leverage a depth renderer to render $\mathbf{P}_{out}$ to depth maps from different viewpoints. These depth maps and those from the category-specific depth image bank are fed to a discriminator, which consists of five 2D Convolutional layers, to discriminate whether they are generated or real samples. It is important to note that the depth image bank is composed of single-view maps, each corresponding one-to-one with the input partial shape, without including samples from additional viewpoints.
Our differentiable rendering pipeline is derived from [68]. Given the intrinsic and extrinsic camera parameters, [68] first transform the point cloud from world to camera coordinates and then project it into the 2D image plane. During rasterization, each point is rendered as a circular splat with a predefined radius in screen space, where the opacity of the splat decreases away from its center. The value of each pixel in the silhouette maps is computed by blending the information from the K nearest points along the z-axis whose splat regions overlap that pixel. For depth maps, each pixel’s value is the z-coordinate of the point that has the smallest z-value among the K nearest points.
However, [68] assigns a fixed and identical radius to each point splat. In fact, the density of 2D points projected by a point cloud with a fixed point number from different viewpoints is different. For example, as shown in Figure 3, when there is a large foreground area of a rendered image, the density of the projected points is small (we call these low-density depth maps). In such cases, a small radius cannot cover all areas of the visible surface. Many “hole pixels” will appear in the rendered depth map. These “hole pixels” are inaccurately represented by the depth values of the points that should have been occluded. In contrast, in scenarios with few foreground pixels, the density of points is large (we call these high-density depth maps). At this time, a small radius suffices to cover the entire visible surface, and thus, the rendered depth map is more realistic and smooth. Therefore, a fixed point radius is inadequate for rendering under different viewpoints.
Intuitively, using a larger point radius for rendering might produce smoother depth maps in both high point density and low point density scenarios. However, in this way, the projected points cover a larger area on the 2D plane, potentially extending the edges of the foreground pixels beyond the original boundaries of the object. This results in a significant overestimation of the object’s size. Moreover, in high-density regions, a large point radius might cause multiple points to overlap excessively. This overlap could obscure finer details, as the rendered depth value for a pixel might be dominated by fewer points, causing the depth map to lose important structural information. Besides, an excessively large point radius may limit the rendering process to updating gradients for only a small subset of points. Therefore, the point radius should dynamically vary with the density of the projected 2D points. The higher the density, the smaller the radius, and vice versa.
To this end, we introduce a density-aware radius estimation (DARE) algorithm to adaptively determine a suitable radius for points from each viewpoint. First, we predefine an initial point radius for the generated points and then render them into a coarse depth map. Note that the predefined point radius should cover the projected area. We then calculate the number of foreground pixels on this depth map, denoted as $A$, and subsequently, the density of the projected points on the depth map is defined as
$d=M/A.$ (6)
Then, we model the relationship between the actual point radius and density with an inversely proportional function and set a coefficient $\eta$ such that
$r=\eta/d,$ (7)
where $r$ denotes the point radius. Note that the parameter $\eta$ can be pre-estimated based on the initial point radii. Specifically, for each category, the density of projected points could be approximated using the average foreground pixel number of all the depth maps in the dataset, as specified in Formula 6. Following this, $\eta$ can be computed according to Formula 7. In this way, the radius of the point cloud dynamically adjusts in response to variations in the density of projected points from various viewpoints, resulting in a smoother rendered depth map.
III-C Loss Functions
To preserve the existing regions in the reconstructed point cloud, we employ a partial matching loss [26], which computes the Unidirectional Chamfer Distance (UCD) from output to input:
$\begin{split}{\rm\texttt{UCD}}(\mathbf{Q}_{1},\mathbf{Q}_{2})=\frac{1}{|\mathbf{Q}_{1}|}\sum_{\mathbf{q}_{1}\in\mathbf{Q}_{1}}\mathop{\texttt{min}}\limits_{\mathbf{q}_{2}\in\mathbf{Q}_{2}}||\mathbf{q}_{1}-\mathbf{q}_{2}||,\end{split}$ (8)
$L_{part}={\rm\texttt{UCD}}(\mathbf{P}_{in},\mathbf{P}_{c})+{\rm\texttt{UCD}}(\mathbf{P}_{in},\mathbf{P}_{out}).$ (9)
Furthermore, we use a rendering loss to supervise the predicted point cloud in 2D space. First, we binarize the depth image $\mathbf{D}_{0}$ to obtain the 2D silhouette ground truth $\mathbf{S}_{0}$ from the input viewpoint. Subsequently, we compute the Mean Squared Error (MSE) loss between $\mathbf{S}_{0}$ and $\mathbf{S}_{out}$. Note that due to the sparsity of the coarse point cloud $\mathbf{P}_{c}$, the foreground pixels of $\mathbf{S}_{c}$ are also sparse and discrete. In contrast, the foreground of $\mathbf{S}_{0}$ is continuous and dense. Therefore, we propose a Masked MSE loss, which only calculates the MSE loss on the foreground pixels of the rendered image. This trick confines the projected points of the rendered image to be accurately positioned.
$\begin{split}L_{rend}&=\frac{1}{H\cdot W}\sum_{i=0}^{H}\sum_{j=0}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{out}(i,j))^{2}\\ &+\frac{1}{H\cdot W}\sum_{i=0}^{H}\sum_{j=0}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{c}(i,j))^{2}\cdot\mathbf{B}(i,j),\end{split}$ (10)
where $\mathbf{B}$ is the binary mask for $\mathbf{S}_{c}$ that indicates whether each pixel of $\mathbf{S}_{c}$ is foreground or not.
To achieve a more uniform point cloud, following [30], we incorporate a density loss $L_{dens}$ to constrain the local density of the point cloud to be as similar as possible.
In addition, the optimization of our generation model is also driven by the adversarial loss proposed by [69], which is defined as:
$L_{gen}=\mathbb{E}_{d_{r}\sim D_{r}}[D(d_{r})-1]^{2}$ (11)
Overall, the generation process of our model is optimized by the generation loss $L_{G}$:
$L_{G}=\alpha_{1}\cdot L_{part}+\alpha_{2}\cdot L_{rend}+\alpha_{3}\cdot L_{dens}+\alpha_{4}\cdot L_{gen},$ (12)
where $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ are the weights for balancing these loss terms. Meanwhile, the discriminator is optimized by $L_{D}$:
$L_{D}=\mathbb{E}_{d_{s}\sim D_{s}}[D(d_{s})-1]^{2}+\mathbb{E}_{d_{r}\sim D_{r}}[D(d_{r})]^{2}$ (13)
[…]
Methods Supervised Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
FoldingNet [40] full 6.41 1.86 9.13 3.19 5.72 12.72 5.42 7.29 6.00
PCN [15] full 5.42 2.22 5.32 3.33 7.25 8.44 5.91 6.48 4.43
TopNet [41] full 4.49 1.65 4.34 3.26 5.25 7.97 5.41 4.49 3.56
AnchorFormer [17] full 3.57 1.06 4.80 2.51 4.23 4.22 4.44 4.64 2.68
PCL2PCL [25] unpaired 16.53 4.68 28.08 6.80 17.01 20.58 14.56 28.88 11.71
KTNet [29] unpaired 6.62 1.41 5.58 4.51 8.95 8.65 8.77 9.86 5.29
USSPA [30] unpaired 14.08 4.01 6.94 6.81 27.36 25.39 12.13 19.31 10.70
ACL-SPC[64] no 15.26 3.38 21.72 5.87 11.38 34.73 13.03 20.24 11.73
P2C [65] no 12.28 1.85 12.73 4.93 12.99 19.43 24.82 14.95 6.57
Ours no 9.08 2.35 10.64 5.92 10.14 14.51 10.15 11.65 7.28
[…]
Methods PCN [15] PCL2PCL [25] ACL-SPC [64] P2C [65] Ours
Metric UCD UHD UCD UHD UCD UHD UCD UHD UCD UHD
chair 5.49 9.19 9.65 10.07 7.74 11.56 7.26 12.21 8.06 9.79
car 2.43 5.22 3.62 5.57 1.54 5.17 1.76 5.02 1.32 4.17
table 11.07 10.80 19.95 13.80 16.55 14.29 9.47 11.03 8.77 10.63
IV Experiments
IV-A Dataset and Evaluation Metric
To comprehensively evaluate our method, we conduct experiments using both synthetic data and real-world data. The synthetic point clouds we used are derived from ShapeNet [70]. We render a single depth map for each object and then back-project it into 3D space to obtain the partial point cloud. The complete point clouds for evaluation are uniformly sampled from the objects’ surfaces. Following PCN [15], we select 30974 objects across 8 categories: airplane, cabinet, car, chair, lamp, couch, table, and watercraft. For each category, we employ 100/150 objects for validation/testing, and the rest is used for training. We set the point numbers of partial input and complete shape to 2048 and 16384, respectively. The evaluation metric for synthetic data is CD, which is widely used in point cloud completion tasks.
For the real-world dataset, we use the data sourced from ScanNet [71] and KITTI [72]. ScanNet contains incomplete objects in indoor scenes, and we extracted 548 chairs and 551 tables from the ScanNet dataset as indoor test data. KITTI provides partial scans of instances in outdoor scenes. We extract 473 partial point clouds from car objects in the Velodyne data, each containing at least 100 points. Due to the lack of complete ground truth, following previous works [15, 25, 27, 16, 32], we adopt the UCD and Unidirectional Hasudorff Distance (UHD) as the evaluation metrics for assessing our model’s performance on the real-world dataset. The UCD is introduced in Formula 8, and the UHD is formulated as:
$\begin{split}{\rm\texttt{UHD}}(\mathbf{Q}_{1},\mathbf{Q}_{2})=\mathop{\texttt{max}}\limits_{q_{1}\in\mathbf{Q}_{1}}\mathop{\texttt{min}}\limits_{q_{2}\in\mathbf{Q}_{2}}\|q_{1}-q_{2}\|_{2}\end{split}$ (14)
IV-B Implementation Details
The point numbers of $\mathbf{P}_{in}$, $\mathbf{P}_{c}$, and $\mathbf{P}_{out}$ are 2048, 2048 and 16384, respectively. We employ neighbor numbers of 16 and 24 for calculating position and curvature encodings, respectively. Initial point radii for the DARE algorithm are set to 0.008 for the plane category and 0.03 for other categories. The loss weights are all set to 1. The initial learning rate is 0.0001, and it decays by 0.5 for every 200 epochs. The training process takes 600 epochs to converge. The experiments are conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
IV-C Comparison with State-of-the-Art Methods
Figure 4: Qualitative results on the synthetic dataset. (a) Partial points. (b) Ground truth. (c) FoldingNet [40]. (d) PCN [15]. (e) AnchorFormer [17]. (f) KTNet [29]. (g) USSPA [30]. (h) ACLSPC [64]. (i) P2C [65]. (j) Ours.
IV-C1 Results on synthetic data
We first conduct experiments on the synthetic dataset and compare the results with other state-of-the-art methods, including fully supervised methods (i.e., FoldingNet [40], PCN [15], TopNet [41], AnchorFormer [17]), unpaired methods (i.e., PCL2PCL [25], KTNet [29], USSPA [30]) and self-supervised methods (i.e., ACL-SPC [64], P2C [65]). The quantitative results presented in Table I suggest that our self-supervised method achieves an impressive average CD of 9.08, significantly outperforming other self-supervised methods and two unpaired approaches (i.e., PCL2PCL and USSPA) by a large margin, and even demonstrating competitive results compared to some fully supervised methods. Regarding the category-specific results, our method significantly outperforms two self-supervised methods across five categories. Specifically, the CD value of MAL-SPC under the lamp category surpasses the results of P2C and ACL-SPC by nearly 5 and 20, respectively. Besides, the CD values of MAL-SPC are lower than those of self-supervised benchmarks by 2-4 under cabinet, couch, and table categories. Furthermore, even with the prior knowledge of complete 3D point clouds, the results of the two unpaired methods, PCL2PCL and USSPA, lag significantly behind our method across multiple categories. In addition, compared to the supervised method FoldingNet, our model achieves a CD difference of less than 3. These results demonstrate the strong performance and effectiveness of our method.
Figure 4 illustrates a comparison of the visualization results between our method and other benchmarks. We can see from Figure 4 that MAL-SPC can infer a more reasonable global shape while better preserving the incomplete parts. For example, the four legs of the input chair point cloud are all thin cylindrical shapes, while other self-supervised or even supervised methods reconstruct chair legs as thick rectangular prisms, whereas MAL-SPC preserves its original shape well. In addition, it is obvious that the noise points generated by our method are significantly fewer than those produced by other methods. Self-supervised benchmarks and most supervised approaches generate a large number of outlier points outside the edges of objects, whereas our method has almost no outlier points.
TABLE III: Comparison of FLOPs and model size.
Methods USSPA [30] ACL-SPC [64] P2C [65] Ours
FLOPs (G) 45.28 0.59 1.79 15.10
Model size (M) 75.06 198.87 427.35 94.79
IV-C2 Results on real-world scans
Figure 5: Qualitative results on the real-world dataset. The first raw represents the input incomplete point cloud, while the second to the last rows correspond to the results of PCN [15], PCL2PCL [25], ACL-SPC [64], P2C [65], and our method, respectively. (a)$\sim$(c) Car. (d)$\sim$(f) Chair. (g)$\sim$(i) Table.
In order to validate the robustness and effectiveness of our method in real-world environments, we evaluate its performance on two challenging datasets: ScanNet and KITTI. As demonstrated in Table II, compared to ACL-SPC and P2C, our MAL-SPC yields the best UCD on all three categories and the best UHD on the car and table categories. Our method only slightly falls behind the second-best approach in terms of UCD on the chair category. Moreover, regarding the results on the car category, MAL-SPC even surpasses the unpaired and supervised methods by a large margin. Figure 5 shows the quantitative results of our model and other state-of-the-art benchmarks. Similar to its superior performance on synthetic data, our method predicts more reasonable shapes and better preserves the original inputs, while generating significantly fewer noise points.
TABLE IV: Quantitative comparison of the effectiveness of each module in our framework. “✓” and “✕” represent the model with or without the component, respectively. “PE” and “CE” denote the position encoding and the curvature encoding, respectively. “MAN” means the multi-view adversarial network, and “DARE” indicates the density-aware radius estimation algorithm. The bold numbers represent the best results.
Component Categories
PE CE MAN DARE Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
11.97 3.89 13.22 7.34 12.83 24.22 11.15 14.43 8.68
10.76 3.85 12.73 6.68 12.40 16.89 11.34 14.12 8.11
9.87 3.44 10.53 5.83 11.75 16.64 10.23 12.73 7.78
9.32 2.25 10.34 6.13 10.97 14.67 10.25 12.52 7.49
9.08 2.35 10.64 5.92 10.14 14.51 10.15 11.65 7.28
IV-C3 Comparison of computational complexity and model size
Table III provides a comprehensive comparison of the computational complexity, measured in Floating Point Operations (FLOPs), and the model size for our proposed method against several state-of-the-art approaches: USSPA, ACL-SPC, and P2C. Our model demonstrates a well-balanced approach in terms of both computational complexity and model size. With 15.10G FLOPs, it is considerably more efficient than USSPA. In terms of model size, our method, with 94.79M parameters, remains manageable and significantly smaller than both ACL-SPC and P2C. This balance makes our model an attractive option for practical applications where both computational resources and memory capacity are constrained.
Figure 6: Visual comparison of reconstructed results with and without the position encodings. With PE, the positions of the predicted point are more accurate, resulting in fewer noisy points. (a) Partial inputs. (b) Point clouds predicted without PE (yellow points) and partial inputs (green points). (c) Point clouds predicted with PE (red points) and partial inputs (green points). (d) Ground truth.
Figure 7: Visual comparison of reconstructed results with and without the curvature encodings. After using CE, the surfaces of the predicted point clouds are smoother, similar to the input incomplete point cloud. (a) Partial inputs. (b) Point clouds predicted without CE. (c) Point clouds predicted with CE. (d) Ground truth.
Figure 8: Comparison of rendered depth maps with a fixed radius and with the DARE algorithm. The first and second rows show the rendering results before and after applying the DARE algorithm, respectively. The DARE algorithm adaptively adjusts the rendering radius based on the density of projected points, enabling smoother foregrounds for low-density depth maps while preventing the expansion of foreground regions in high-density depth maps. (a)$\sim$(d) High-density depth map. (e)$\sim$(h) Low-density depth map.
IV-D Ablation Study
IV-D1 Effects of on each module
To analyze the effect of each innovative component in our framework, including position encoding, curvature encoding, multi-view adversarial network, and density-aware radius estimation strategy, we conducted ablation experiments on synthetic data. As Table IV shows, the inclusion of all modules (PE, CE, MAN, and DARE) yields the best average performance, achieving the lowest average error of 9.08. In particular, adding only PE results in a noticeable reduction in the average error to 10.76, especially a substantial gain in the lamp category (from 24.22 to 16.89). We also compared the visualization results with and without using PE to visually observe the impact of PE on the reconstruction process. As Figure 6 shows, without PE, the model predicts some inaccurate points that significantly deviate from the partial input and the ground truth. However, after incorporating PE, these inaccurate points are noticeably reduced. We can conclude from both the quantitative and qualitative results that PE is able to effectively utilize the positional relationships among partial points to infer more accurate point positions.
Additionally, it is evident from Table IV that incorporating both PE and CE leads to a further reduction in the average error to 9.87. CE appears to complement PE by providing additional curvature information. The combination significantly improves performance in categories like table, couch, and cabinet. Moreover, as observed from the visual results in Figure 7, after embedding CE, when the surface of the incomplete input point cloud is compact and smooth, the corresponding reconstructed surface is also significantly smoother. These demonstrate that CE can effectively learn the curvature features of the partial input and leverage them to reconstruct more realistic surfaces.
In addition to PE and CE, the inclusion of MAN further enhances the performance as demonstrated in Table IV, with an average improvement of 0.55. Note that the density-aware radius estimation algorithm also contributes positively to the model’s performance, with significant improvements in categories like chair and table.
To more intuitively demonstrate the effect of the DARE algorithm, we compared the depth maps rendered with a fixed radius and with the DARE algorithm. As shown in Figure 8, when a fixed radius is used for rendering, low-density depth maps appear smoother, but high-density depth maps often contain many holes. However, after applying the DARE algorithm, the foreground of high-density depth maps becomes smoother, while the edges of foreground pixels in low-density depth maps do not expand. This proves that the proposed DARE algorithm can adaptively select an appropriate point radius based on the density of projected points, thereby rendering higher-quality depth maps.
IV-D2 Ablation on supervisory signals
Our solution employs two types of direct supervisory signals: an incomplete 3D point cloud and a 2D depth map. Although these signals have different data representations, they both originate from the same source, i.e., the incomplete point cloud is derived by back-projecting the depth map into 3D space. We do not incorporate 2D supervision from other viewpoints.
To evaluate the impact of these two supervisory signals on the model’s performance, we conducted ablation studies on the synthetic dataset. We used our complete model and performed experiments by separately removing $L_{part}$ and $L_{rend}$. The results show that non-converge occurs in the absence of either $L_{part}$ or $L_{rend}$. For the case of lacking $L_{rend}$, the CD values achieved on the cabinet, car, couch, and table categories are 36.56, 18.19, 19.68, and 74.83, respectively, while convergence is not achieved on other categories. On the other hand, when $L_{part}$ is removed, CD values on cabinet, chair, car, couch, table, and watercraft categories are 29.13, 21.12, 7.36, 12.00, 12.53, and 9.78, respectively, while convergence failed for plane and lamp categories. We believe that the reason for the non-convergence is twofold. First, the training of GAN itself is inherently unstable [62, 69]. Second, the inclusion of DARE makes training more challenging. This is because DARE requires a relatively accurate initial depth map rendered from the predicted results before estimating a more precise point radius. If the initial depth map has significant deviations, the estimated point radius will also deviate accordingly, leading to significant errors in the second-round rendering. Consequently, the GAN module might collapse, ultimately preventing the model from converging. The lack of either $L_{rend}$ or $L_{part}$, which provides direct supervision, increases the difficulty for the model to predict accurate shapes, contributing to the non-convergence. However, based on the results from the converged categories, $L_{rend}$ exhibits stronger supervisory capabilities compared to $L_{part}$.
V Conclusion
In the scenario of self-supervised point cloud completion, it is critical to fully utilize the partial observations in the training data as priors due to the lack of complete shapes as supervision. To this end, we have presented a novel approach that effectively exploits the objects’ intrinsic object-level and category-specific geometric similarity to infer complete shapes without the need for multi-view partial scans and masked operation. Our method demonstrates strong capabilities in both inferring complete shapes and preserving incomplete input, resulting in reconstructed shapes that are both reasonable and accurate, with minimal noise. Quantitatively and visually, our approach significantly surpasses current state-of-the-art self-supervised methods, as well as a portion of unpaired methods, by a large margin. In future work, we aim to extend our work to large-scale indoor and outdoor scene completion tasks. We hope this work will inspire future research in 3D object and scene completion, reconstruction, and understanding tasks.